# Beyond the Permutation Symmetry of Transformers: The Role of Rotation for Model Fusion

### v1

@@ 1 Introduction
Parameter space symmetry is an intriguing property of neural networks that has garnered increasing attention in recent years (Du et al., 2018; Armenta & Jodoin, 2021; Kunin et al., 2021; Simsek et al., 2021; Entezari et al., 2022; Grigsby et al., 2023). One of the most studied forms of parameter space symmetry is permutation symmetry (Ainsworth et al., 2023; Entezari et al., 2022). For instance, in a two-layer MLP, permuting the rows of the weight matrix in the first layer and applying the corresponding inverse permutation to the second layer results in a functionally equivalent model, i.e., the outputs of the original and permuted models remain identical for any given input (Ainsworth et al., 2023). All functionally equivalent models corresponding to weight permutations form an equivalence set, which provides theoretical insights into neural network optimization, such as the linear mode connectivity of loss landscapes (Entezari et al., 2022; Zhou et al., 2023). In addition, permutation symmetry has also proven helpful in advancing neural network applications, such as model fusion (Singh & Jaggi, 2020; Ainsworth et al., 2023) and optimization (Zhao et al., 2024).
Although parameter space symmetry has been extensively studied in classical neural network architectures, such as MLPs and CNNs, the understanding of its application in transformers (Vaswani et al., 2017) remains limited. Transformers have seen rapid advancements in recent years, achieving remarkable success in a wide range of applications (Yun et al., 2019; Lewis et al., 2020; Raffel et al., 2020; Clark et al., 2020; Liu et al., 2021; Dosovitskiy et al., 2021; Zhou et al., 2021; He et al., 2024; Zhu et al., 2024; Zheng et al., 2024). The transformer architecture is built upon two primary submodules: feedforward networks and self-attention layers. The feedforward network, which is structurally similar to MLPs, naturally inherits the permutation symmetry that has been extensively studied in the existing literature. Self-attention layers, on the other hand, involve a unique attention mechanism powered by matrix products of queries, keys, and values, which introduce additional potentials for symmetry beyond permutations. Permutation symmetries limit the equivalence set of neural networks to discrete operations, which aligns well with MLPs due to their element-wise activations (e.g., ReLU (Glorot et al., 2011)). In contrast, the continuous nature of the matrix operations in self-attention layers necessitates more flexible operations to fully characterize their equivalence set.
@@ 4.2 Parameter Matching
We leave a detailed proof of Theorem 4.1 in the appendix. According to Theorem 4.1, we can obtain the algorithm to match two self-attention layers shown in Algorithm 1. Algorithm 1 can be seen as an adaptation of the Kabsch algorithm (Kabsch, 1976; Umeyama, 1991). Without loss of generality, we let the parameters in the second ($k=2$) self-attention layer be the anchor and conduct rotation for the first layer ($k=1$). The denseness of rotation symmetry helps reduce the distance between the end models after parameter matching as Algorithm 1 shows.
Scaling Matching.
We find that the rescaling symmetry (Neyshabur et al., 2015; Meng et al., 2019) can be leveraged jointly with our proposed rotation symmetry in the parameter matching algorithm. For instance, for the (simplified) Q-K product $\bm{W}_{Q}\bm{W}_{K}$, we can find a rescaling operation that preserves the functionality of the Q-K product $a\bm{W}_{Q}\cdot\frac{1}{a}\bm{W}_{K}$ where $a\neq 0$ is a real number. By adding a scalar variable to each parameter matrix, we can formulate the objective of rescaling symmetry as follows (taking the Q-K product as an example).
@@
$+\left\|\bm{W}_{Q_{1}}^{\prime}/a-\bm{W}_{Q_{2}}^{\prime}\right\|^{2}+\left\|\bm{b}_{Q_{1}}^{\prime}/a-\bm{b}_{Q_{2}}^{\prime}\right\|^{2},$
where $a$ is the rescaling variable and denotes the parameters after rotation symmetry-based matching. Here, we still set model 2 as the anchor model and conduct the rescaling operation to model 1 to align with model 2. We can easily solve the optimality condition of Equation 11 as
@@
We next provide a brief analysis of the complexity of our proposed parameter matching algorithm. We let the hidden dimension of the target transformer be $d$, the layer of the target transformer be $L$, and the number of attention heads be $H$. For the parameter matching of feedforward networks, the linear assignment problem can be solved in $O(d^{3})$ by the Hungarian algorithm (Kuhn, 1955; Martello & Toth, 1987). Additionally, to solve the optimization problem in Equation 8, Algorithm 1 requires the complexity of $O(d^{3})$ for eigendecomposition. Hence, the complexity of our proposed full parameter matching algorithm of a transformer is $O(d^{3}LH)$, similar to the complexity of the feedforward. It is worth noting that the complexity can be further reduced in two ways. The first way is to match a subset of layers instead of all. In this way, the complexity of our proposed parameter matching algorithm is $O(d^{3}L_{s}H)$ where $L_{s}$ denotes the number of selected layers. The second way is to match each unit module (a single feedforward network or a single attention layer) in parallel. The decoupling of matching different modules makes it easy to implement multiprocessing and the overall complexity becomes $O(d^{3}LH/p)$ where $p$ is the number of processes in parallel.
@@
Table 2: Experimental results of ViT merging over image classification task. OT represents OT-acts-EMD.
Simple Fisher Regmean OT
w/o matching 7.60 17.96 14.24 32.08
w/ matching 10.22 18.61 15.31 32.50
w/ scaling matching 10.19 18.58 15.35 32.53
@@ 5 Experiments
In this section, we conduct experiments to evaluate the effectiveness of our proposed parameter matching algorithm based on rotation symmetry. Specifically, we aim to answer the following research questions: RQ1: Can our proposed parameter matching algorithm enhance the performance of model fusion for transformer-based models? RQ2: How does rotation symmetry contribute to parameter matching for self-attention layers? RQ3: As a plugin module, does our algorithm introduce significant additional computational overhead? RQ4: Is matching all transformer layers equally important? Can we improve efficiency without compromising utility by matching only a subset of layers?
@@ 5.1 Experimental Settings
In the experiments, each pretrained language model is fine-tuned for 20 epochs on each dataset individually. We set the learning rate at 1e-5, the batch size at 16, and the warmup ratio at 0.06 for each model. After fine-tuning, we perform parameter matching and model merging, considering both in-domain (pairwise fine-tuned models) and out-of-domain (grouped models) experiments. Notably, we match only the parameters within the attention layers, while the classifier module is directly copied from the model fine-tuned on the corresponding downstream task. For additional details on datasets and baseline methods, please refer to Appendix D.
(a) Single-Layer Matching
(b) Tail-Layers Matching
Figure 4: Evaluation loss of matching a subset of layers.
@@ 5.2 Performance of Model Fusion
We employ OT-acts-EMD from (Imfeld et al., 2024) as an additional baseline. Two pretrained models are merged, and we evaluate their performance on the CIFAR-10 (Krizhevsky et al., 2009) dataset. Table 2 presents the classification accuracy of the merged models. To improve the flexibility of parameter matching, we match each layer separately before conducting model fusion. The merged model achieving the highest validation performance is selected as the final test model.
From the results in Table 1 and Table 2, we make the following observations: (1). Our parameter matching algorithm consistently improves the performance of different model fusion methods. The reason is that parameter matching helps align distant models, bringing them closer in parameter space. As a result, the merged model is more likely to approach an overall minimum. (2). Compared to RoBERTa-base, our method brings larger improvement for DeBERTa-large, suggesting that larger models benefit more from parameter matching. We explain this as larger models have more parameters that can be aligned, and in high-dimensional spaces, models tend to be more spread out. Subsequently, parameter matching helps bridge this gap more effectively. (3). Among all the fusion methods, simple fusion shows the largest improvement after parameter matching. This is likely because simple fusion directly averages model weights, so reducing the distance between model parameters brings a clear benefit. In contrast, advanced fusion methods work by aligning outputs of the end models based on input data, which is equivalent to weighted averaging. Unlike direct averaging, these methods benefit from parameter matching in a more implicit way.
@@ 5.3 Ablation Study
To answer RQ2, we conduct ablation studies to evaluate the impact of rotation symmetry in parameter matching. Specifically, match the two pretrained ViT models utilized in the previous section. To assess the impact of rotation symmetry, we measure the Euclidean distance between the parameters in the matched model and the anchor model. Instead of aggregating distances at the layer or module level, we analyze the entire model holistically to provide a more comprehensive assessment of alignment. For comparison, we also include two external baselines: Git-Rebasin (Ainsworth et al., 2023) and OT-Fusion (Imfeld et al., 2024). Following the original paper, we apply OT-Fusion exclusively to the feed-forward networks. The results in Figure 3 show that our rotation symmetry-based parameter matching algorithm consistently reduces the distance between end models more effectively than permutation symmetry-based methods. The advantage of rotation symmetry lies in its ability to operate in a continuous space, allowing for smoother and more effective parameter alignment in self-attention layers. Additionally, incorporating rescaling symmetry further refines parameter alignment, leading to a greater distance reduction.
@@ 5.4 Complexity Study
To answer RQ3, we evaluate the computational overhead introduced by our parameter matching algorithm. Specifically, we measure the average runtime for fine-tuning (per dataset), matching (per model pair), and merging methods (per model pair) from the main experiments, as shown in Table 3. For ViTs, we only use pre-trained models following (Imfeld et al., 2024) without fine-tuning. The results show that our parameter matching module incurs an overhead of less than 5% compared to Fisher / RegMean merging across all models, with negligible impact on overall complexity.
@@ 5.5 Matching a Subset of Layers
To answer RQ4, we investigate the effect of matching different subsets of layers. We fine-tune DeBERTa models on CoLA and STS-B from the GLUE benchmark and evaluate the performance of matched models under different subsets of layers. Specifically, we cross-validate the importance of each attention layer in matching through two settings:
@@ 6 Related Work
Parameter space symmetry refers to a set of models with different parameter values but functionally equivalent. This concept has been extensively studied in the context of deep neural networks, as it plays a crucial role in understanding model behavior and training dynamics. Examples of parameter space symmetries include rescaling symmetry (Neyshabur et al., 2015; Badrinarayanan et al., 2015; Du et al., 2018; Meng et al., 2019), scaling symmetry (Kunin et al., 2021), and translation symmetry (Kunin et al., 2021). These symmetries have been identified in conventional deep neural networks to provide deeper insights into training dynamics and to accelerate the optimization process (Zhao et al., 2022, 2023, 2024). Another important type of parameter space symmetry is permutation symmetry, which has been shown to closely relate to the manifold of global minima and critical points (Fukumizu & Amari, 2000; Brea et al., 2019; Simsek et al., 2021; Benton et al., 2021; Entezari et al., 2022; Ainsworth et al., 2023). The permutation symmetry can also be used to align (match) the outputs or model parameters of different end models with the same architecture (Singh & Jaggi, 2020; Wang et al., 2020; Ainsworth et al., 2023; Imfeld et al., 2024). Some concurrent works (Liu, 2024; Tran et al., 2024) investigate similar forms of rotation symmetry in neural networks. Liu (2024) shows that the mirror symmetry leads to low-rankness. Meanwhile, Tran et al. (2024) leverages the rotation symmetry to construct transformer-based neural functional networks. In comparison, our study focuses on the role of rotation symmetry in model fusion and proposes a theoretically optimal parameter matching approach based on the properties of rotation symmetries.
@@
The goal of model fusion (Li et al., 2023) is to merge multiple available end models (with the same architecture) to obtain a stronger model. The scenarios of model fusion can be flexible. When training on the same dataset, model fusion can be used to improve the model utility or generalization by merging models trained with different configurations or in different stages (Izmailov et al., 2018; Gupta et al., 2020; Cha et al., 2021; Wortsman et al., 2022; Rame et al., 2022; Arpit et al., 2022). As a representative method in this setting, ModelSoup (Wortsman et al., 2022) greedily averages the models fine-tuned with different hyperparameter configurations to improve the utility and robustness of the model. In addition, when training on different datasets or tasks, model fusion can be used to improve out-of-domain generalization or multitasking of the model (Matena & Raffel, 2022; Choshen et al., 2022; Li et al., 2022; Jin et al., 2023), especially for language models. A state-of-the-art merging algorithm, RegMean (Jin et al., 2023), successfully merges language models fine-tuned over different tasks and improves the model’s out-of-distribution generalization. Moreover, model fusion plays a pivotal role in federated learning (Konečnỳ et al., 2016; McMahan et al., 2017; Wang et al., 2020) when the local updates are collected to make a global update. FedAvg (McMahan et al., 2017) is a classical merging algorithm that directly computes the average of the local models as the updated global model. Recent studies propose to incorporate the permutation symmetry to align the neurons of different end models (Wang et al., 2020; Singh & Jaggi, 2020; Ainsworth et al., 2023). However, these methods fail to achieve a desirable performance when tackling transformer-based models (Jin et al., 2023).

### v2

@@ 1 Introduction
Parameter space symmetry is an intriguing property of neural networks that has garnered increasing attention in recent years (Du et al., 2018; Armenta & Jodoin, 2021; Kunin et al., 2021; Simsek et al., 2021; Entezari et al., 2022; Grigsby et al., 2023; Lim et al., 2024b; Ziyin et al., 2024, 2025). One of the most studied forms of parameter space symmetry is permutation symmetry (Ainsworth et al., 2023; Entezari et al., 2022). For instance, in a two-layer MLP, permuting the rows of the weight matrix in the first layer and applying the corresponding inverse permutation to the second layer results in a functionally equivalent model, i.e., the outputs of the original and permuted models remain identical for any given input (Ainsworth et al., 2023). All functionally equivalent models corresponding to weight permutations form an equivalence set, which provides theoretical insights into neural network optimization, such as the linear mode connectivity of loss landscapes (Entezari et al., 2022; Zhou et al., 2023c; Ferbach et al., 2024). In addition, permutation symmetry has also proven helpful in advancing neural network applications, such as model fusion (Singh & Jaggi, 2020; Ainsworth et al., 2023) and optimization (Zhao et al., 2024; Zamir et al., 2025).
Although parameter space symmetry has been extensively studied in classical neural network architectures, such as MLPs and CNNs, the understanding of its application in transformers (Vaswani et al., 2017) remains limited. Transformers have seen rapid advancements in recent years, achieving remarkable success in a wide range of applications (Yun et al., 2019; Lewis et al., 2020; Raffel et al., 2020; Clark et al., 2020; Zhou et al., 2021; He et al., 2024; Zhu et al., 2024; Zheng et al., 2024; Zhang et al., 2025; Wei & Liu, 2025). The transformer architecture is built upon two primary submodules: feedforward networks and self-attention layers. The feedforward network, which is structurally similar to MLPs, naturally inherits the permutation symmetry that has been extensively studied in the existing literature. Self-attention layers, on the other hand, involve a unique attention mechanism powered by matrix products of queries, keys, and values, which introduce additional potentials for symmetry beyond permutations. Permutation symmetries limit the equivalence set of neural networks to discrete operations, which aligns well with MLPs due to their element-wise activations (e.g., ReLU (Glorot et al., 2011)). In contrast, the continuous nature of the matrix operations in self-attention layers necessitates more flexible operations to fully characterize their equivalence set.
@@ 4.2 Parameter Matching
We leave a detailed proof of Theorem 4.1 in Appendix A. According to Theorem 4.1, we can obtain the algorithm to match two self-attention layers shown in Algorithm 1. Algorithm 1 can be seen as an adaptation of the Kabsch algorithm (Kabsch, 1976; Umeyama, 1991). Without loss of generality, we let the parameters in the second ($k=2$) self-attention layer be the anchor and conduct rotation for the first layer ($k=1$). The denseness of rotation symmetry helps reduce the distance between the end models after parameter matching as Algorithm 1 shows.
Algorithm 1 Matching two self-attention layers.
Rescaling Matching.
We find that the rescaling symmetry (Neyshabur et al., 2015; Meng et al., 2019; Godfrey et al., 2022; Kalogeropoulos et al., 2024) can be integrated with our proposed rotation symmetry in the parameter matching algorithm. For instance, for the (simplified) Q-K product $\bm{W}_{Q}\bm{W}_{K}$, we can find a rescaling operation that preserves the functionality of the Q-K product $a\bm{W}_{Q}\cdot\frac{1}{a}\bm{W}_{K}$ where $a\neq 0$ is a real number. By adding a scalar variable to each parameter matrix, we can formulate the objective of rescaling symmetry as follows (taking the Q-K product as an example).
@@
$+\left\|\bm{W}_{K_{1}}^{\prime}/a-\bm{W}_{K_{2}}^{\prime}\right\|^{2}+\left\|\bm{b}_{K_{1}}^{\prime}/a-\bm{b}_{K_{2}}^{\prime}\right\|^{2},$
where $a$ is the rescaling variable and denotes the parameters after rotation symmetry-based matching. Here, we still set model 2 as the anchor model and conduct the rescaling operation on model 1 to align with model 2. We can easily solve the optimality condition of Equation 11 as
@@
We next provide a brief analysis of the complexity of our proposed parameter matching algorithm. We let the hidden dimension of the target transformer be $d$, the layer of the target transformer be $L$, and the number of attention heads be $H$. For the parameter matching of feedforward networks, the linear assignment problem can be solved in $O(d^{3})$ by the Hungarian algorithm (Kuhn, 1955; Martello & Toth, 1987). Additionally, to solve the optimization problem in Equation 8, Algorithm 1 requires the complexity of $O(d^{3})$ for eigendecomposition. Hence, the complexity of our proposed full parameter matching algorithm of a transformer is $O(d^{3}LH)$, similar to the complexity of the feedforward. It is worth noting that the complexity can be further reduced in two ways. The first way is to match a subset of layers instead of all. In this way, the complexity of our proposed parameter matching algorithm is $O(d^{3}L_{s}H)$ where $L_{s}$ denotes the number of selected layers. The second way is to match each unit module (a single feedforward network or a single attention layer) in parallel. The decoupling of matching different modules makes it easy to implement multiprocessing. Consequently, the overall complexity becomes to $O(d^{3}LH/p)$ where $p$ is the number of processes in parallel.
@@
Table 2: Experimental results of ViT merging over the image classification task.
w/o match w/o rescale match
Simple 7.60 10.22 10.19
Fisher 17.96 18.61 18.58
Regmean 14.24 15.31 15.35
OT-ACTS-EMD 32.08 32.50 32.53
OT-ACTS 61.15 61.23 61.25
OT-WTS 57.11 57.16 57.17
Figure 3: Ablation Study of ViT merging over the image classification task. “ATTN” is short for “attention”. We compare our matching algorithm with its three variants (w/o ATTN/FFN/rescaling) and the original performance (w/o match) on all six merging baselines.
@@ 5 Experiments
In this section, we conduct experiments to evaluate the effectiveness of our proposed parameter matching algorithm based on rotation symmetry. Specifically, we aim to answer the following research questions: RQ1: Can our proposed parameter matching algorithm enhance the performance of model fusion for transformer-based models? RQ2: How does rotation symmetry contribute to parameter matching for self-attention layers? RQ3: How does parameter matching influence the loss landscape between independently trained models? RQ4: As a plugin module, does our algorithm introduce significant additional computational overhead? RQ5: Is matching all transformer layers equally important? Can we improve efficiency without compromising utility by matching only a subset of layers?
@@ 5.1 Experimental Settings
In the experiments, each pretrained language model is fine-tuned for 20 epochs on each dataset individually. We set the learning rate at 1e-5, the batch size at 16, and the warmup ratio at 0.06 for each model. After fine-tuning, we perform parameter matching and model merging, considering both in-domain (pairwise fine-tuned models) and out-of-domain (grouped models) experiments. Notably, we match only the parameters within the attention layers, while the classifier module is directly copied from the model fine-tuned on the corresponding downstream task. For additional details on datasets and baseline methods, please refer to Appendix C.
@@ 5.2 Performance of Model Fusion
We employ three OT (optimal transportation)-based methods (OT-ACTS-EMD, OT-ACTS, and OT-WTS) from (Imfeld et al., 2024) as additional baselines. Two pretrained models are merged, and we evaluate their performance on the CIFAR-10 (Krizhevsky et al., 2009) dataset. Table 2 presents the classification accuracy of the merged models. To improve the flexibility of parameter matching, we match each layer separately before conducting model fusion. The merged model achieving the highest validation performance is selected as the final test model.
From the results in Table 1 and Table 2, we make the following observations: (1) Our parameter matching algorithm consistently improves the performance of different model fusion methods. The reason is that parameter matching helps align distant models, bringing them closer in parameter space. As a result, the merged model is more likely to approach an overall minimum. (2) Compared to RoBERTa-base, our method brings larger improvement for DeBERTa-large, suggesting that larger models benefit more from parameter matching. We explain this as larger models have more parameters that can be aligned, and in high-dimensional spaces, models tend to be more spread out. Subsequently, parameter matching helps bridge this gap more effectively. (3) Among all the fusion methods, simple fusion shows the largest improvement after parameter matching. This is likely because simple fusion directly averages model weights, so reducing the distance between model parameters brings a clear benefit. In contrast, advanced fusion methods work by aligning outputs of the end models based on input data, which is equivalent to weighted averaging. Unlike direct averaging, these methods benefit from parameter matching in a more implicit way.
@@ 5.3 Ablation Study
To answer RQ2, we conduct ablation studies to evaluate the contributions of rotation symmetry in matching different model components. These experiments are performed on the same pair of pretrained ViT models used in the previous section. We design three ablation settings: (1) w/o ATTN: Matches only the FFN modules, excluding self-attention modules. (2) w/o FFN: Matches only the self-attention modules, excluding FFN modules. (3) w/o rescaling: Disables the rescaling symmetry in our algorithm. Ablation results in Figure 3 demonstrate the following observations: (1) In Fisher and RegMean, attention matching contributes more significantly to the fusion process. (2) In contrast, OTFusion benefits more from FFN matching. These distinctions demonstrate the nuanced interactions between the fusion mechanisms and corresponding component alignment strategies. (3) w/o rescaling results in performance degradation across most settings, validating the effectiveness of rescaling symmetry integrated with rotation symmetry.
To further quantify the influence of rotation symmetry, we measure the Euclidean distance between the matched model and the anchor model in the parameter space. Instead of aggregating distances at the layer or module level, we analyze the entire model holistically to provide a more comprehensive assessment of alignment. We also compare our method against two external baselines: Git Re-Basin (Ainsworth et al., 2023) and OT-Fusion (Imfeld et al., 2024). Following the original paper, OT-Fusion is applied only to FFN modules. The results in Figure 4 indicate that our rotation symmetry-based parameter matching algorithm consistently reduces the distance between end models more effectively than permutation symmetry-based methods. This improvement can be attributed to the continuous nature of rotation symmetry, which allows for smoother and more precise alignment, especially in self-attention layers. Furthermore, the integration of rescaling symmetry leads to additional distance reductions of end models, further enhancing parameter alignment.
Figure 5: Loss landscapes and barriers between the two pretrained ViT models under four distinct matching settings. “LB” is short for “Loss Barrier”.
@@
5.4 Loss Landscape Study
To address RQ3, we examine the loss landscape between the two pretrained ViT models introduced in Section 5.2, both with and without parameter matching. Specifically, we first analyze the cross entropy loss values $\mathcal{L}(\theta(\alpha))$ along the linear interpolation path between the end model parameters $\theta(\alpha)=\alpha\cdot\theta_{A}+(1-\alpha)\cdot\theta_{B},\quad\alpha\in[0,1]$, where $\theta_{A}$ and $\theta_{B}$ denote the end models. Next, we separately apply attention-only matching (w/o FFN), FFN-only matching (w/o ATTN), and complete matching (w/ match) to $\theta_{B}$ before constructing the same linear interpolation path with respect to $\theta_{A}$. The results in Figure 5 give rise to the following observations: (1) All model pairs exhibit positive loss barriers, i.e., the range of loss value along the interpolation path, highlighting the strong inherent non-convexity of ViTs (Park & Kim, 2022). (2) Complete matching yields the lowest loss barrier, indicating that aligning parameters through rotation symmetry leads to smoother connectivity between model pairs. (3) While FFN-only matching (w/o ATTN) has been proven effective for MLPs and CNNs (Ainsworth et al., 2023), its effectiveness remains limited for Transformers. Their intricate loss landscape brings additional challenges for parameter matching.
@@ 5.5 Complexity Study
To answer RQ4, we evaluate the computational overhead introduced by our parameter matching algorithm. Specifically, we measure the average runtime for fine-tuning (per dataset), matching (per model pair), and merging methods (per model pair) from the main experiments, as shown in Table 3. For ViTs, we only use pre-trained models following (Imfeld et al., 2024) without fine-tuning. The results show that our parameter matching module incurs an overhead of less than 5% compared to Fisher / RegMean merging across all models, with negligible impact on overall complexity.
(a) Single-Layer Matching
(b) Tail-Layers Matching
Figure 6: Evaluation loss of matching a subset of layers.
@@ 5.6 Matching a Subset of Layers
To answer RQ5, we investigate the effect of matching different subsets of layers. We fine-tune DeBERTa models on CoLA and STS-B from the GLUE benchmark and evaluate the performance of matched models under different subsets of layers. Specifically, we cross-validate the importance of each attention layer in matching through two settings:
@@ 6 Related Work
Parameter space symmetry refers to a set of models with different parameter values but functionally equivalent. This concept has been extensively studied in the context of deep neural networks, as it plays a crucial role in understanding model behavior and training dynamics. Examples of parameter space symmetries include rescaling symmetry (Neyshabur et al., 2015; Badrinarayanan et al., 2015; Du et al., 2018; Meng et al., 2019), scaling symmetry (Kunin et al., 2021), and translation symmetry (Kunin et al., 2021). These symmetries have been identified in conventional deep neural networks to provide deeper insights into training dynamics and to accelerate the optimization process (Zhao et al., 2022, 2023, 2024). Another important type of parameter space symmetry is permutation symmetry, which has been shown to closely relate to the manifold of global minima and critical points (Fukumizu & Amari, 2000; Brea et al., 2019; Simsek et al., 2021; Benton et al., 2021; Entezari et al., 2022; Ainsworth et al., 2023). The permutation symmetry can also be used to align (match) the outputs or model parameters of different end models with the same architecture (Singh & Jaggi, 2020; Wang et al., 2020; Ainsworth et al., 2023; Peña et al., 2023; Imfeld et al., 2024; Navon et al., 2024). Complementary to parameter space symmetry, recent advances in neural functionals and metanetworks (Navon et al., 2023; Zhou et al., 2023b, a; Lim et al., 2024a; Tran et al., 2024) explore permutation-equivariant functionals that operate directly on model weights across diverse architectures. Some concurrent works (Liu, 2024; Tran et al., 2024) investigate similar forms of rotation symmetry in neural networks. Liu (2024) shows that the mirror symmetry leads to low-rankness. Meanwhile, Tran et al. (2024) leverages the rotation symmetry to construct transformer-based neural functional networks. In comparison, our study focuses on the role of rotation symmetry in model fusion and proposes a theoretically optimal parameter matching approach based on the properties of rotation symmetries.
@@
The goal of model fusion (Li et al., 2023) is to merge multiple available end models (with the same architecture) to obtain a stronger model. The scenarios of model fusion can be flexible. When training on the same dataset, model fusion can be used to improve the model utility or generalization by merging models trained with different configurations or in different stages (Izmailov et al., 2018; Gupta et al., 2020; Cha et al., 2021; Wortsman et al., 2022; Rame et al., 2022; Arpit et al., 2022; Huang et al., 2024; Yadav et al., 2024; Hammoud et al., 2024). As a representative method in this setting, ModelSoup (Wortsman et al., 2022) greedily averages the models fine-tuned with different hyperparameter configurations to improve the utility and robustness of the model. In addition, when training on different datasets or tasks, model fusion can be used to improve out-of-domain generalization or multitasking of the model (Matena & Raffel, 2022; Choshen et al., 2022; Li et al., 2022; Jin et al., 2023; Zhou et al., 2024), especially for language models. A state-of-the-art merging algorithm, RegMean (Jin et al., 2023), successfully merges language models fine-tuned over different tasks and improves the model’s out-of-distribution generalization. Moreover, model fusion plays a pivotal role in federated learning (Konečnỳ et al., 2016; McMahan et al., 2017; Wang et al., 2020) when the local updates are collected to make a global update. FedAvg (McMahan et al., 2017) is a classical merging algorithm that directly computes the average of the local models as the updated global model. Recent studies propose to incorporate the permutation symmetry to align the neurons of different end models (Wang et al., 2020; Singh & Jaggi, 2020; Ainsworth et al., 2023). However, these methods fail to achieve a desirable performance when tackling transformer-based models (Jin et al., 2023).